# On Erlang mixture approximations for differential equations with distributed time delays

### v1

@@ 6.2 Identify kernel parameters
where the quadrature points are $t_{k}=k\Delta t$ for $k=0,\ldots,N-1$ and $a_{\min}\in\mathbb{R}_{>0}$ is an arbitrary lower bound, which ensures that only positive values of $a$ are feasible. Furthermore, the objective function involves the regular kernel, $\alpha:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}$, and the $M$’th order Erlang mixture kernel, $\hat{\alpha}:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}$, with coefficients $\{c_{m}\}_{m=0}^{M}$ and rate parameter $a$, evaluated at the quadrature points.
@@ 6.4 Implementation
We implement the evaluation of the objective function in the NLP (95) and its first- and second-order derivatives in C, and we use Matlab’s MEX interface to call the C routines from Matlab. Furthermore, we use Matlab’s fmincon to approximate the solution to the NLP, and we provide analytical first- and second-order derivatives. We exploit that $\alpha_{m}(t)=(at/m)\alpha_{m-1}(t)$ for $m=1,\ldots,M$ where $\alpha_{0}(t)=ae^{-at}$ is less prone to rounding errors than a straightforward evaluation of the expression (10) in the definition, and we use either double or long double precision when evaluating the Erlang kernels, the Erlang mixture kernels, and their derivatives. We simulate the approximate system of ODEs (61) using either Matlab’s ode45 or ode15s, and depending on the order of the Erlang mixture kernel, we implement the matrices $A$, $B$, and $C$ as either dense or sparse. Furthermore, we provide the analytical Jacobian of the right-hand side function of the approximate system of ODEs (61) to the simulator. We implement the numerical method for non-stiff DDEs described in Appendix B.1 in both Matlab and C (with a MEX interface). The numerical method from Appendix B.2 is implemented in Matlab, we use fsolve to solve the residual equations, and we provide the analytical Jacobian. In all three implementations, the right-hand side function, $f$, the delay function, $h$, and the kernel, $\alpha$, are evaluated in Matlab. Finally, in Section 8, we carry out parallel simulations and kernel approximations using a high-performance computing cluster [14].
@@ 7 Numerical test of the accuracy
In this section, we construct a DDE in the form (2b)–(3) with a known solution and use it to investigate the accuracy of the Erlang mixture approximation and the numerical solution of the approximate system of ODEs (61). Furthermore, we assess the convergence rate of the two numerical methods for solving initial value problems in the form (2)–(3) described in Appendix B.
@@
(104) $E_{x}$ $=\sum_{n=0}^{K_{x}-1}(\hat{x}(t_{n+1})-x^{*}(t_{n+1}))^{2}\Delta t,$ $E_{\alpha}$ $=\sum_{k=0}^{K_{\alpha}-1}(\hat{\alpha}(t_{k})-\alpha(t_{k}))^{2}\Delta t.$
@@
Here, $\hat{x}$ and $\hat{\alpha}$ denote the approximations. Note that the kernel error, $E_{\alpha}$, is analogous to the objective function in (95a) but may contain a different number of terms. We have used the model parameter values $\sigma=4$ mo-1 and $\kappa=1$ mo and the dilation parameter $\gamma=10$ mo, and the initial and final times are $t_{0}=0$ mo and $t_{f}=24$ mo. When we determine the Erlang mixture approximation of a given order, $M$, we use $N=10^{2}$ points in the objective function in (95a). Furthermore, we use a threshold of $\epsilon=10^{-14}$ and a convergence tolerance of $10^{-15}$ to determine the approximation interval length, $t_{h}$. Finally, we solve the NLP (95) using Matlab’s fmincon with an optimality tolerance of $10^{-10}$ and a step size tolerance below machine precision, and we use double precision in the involved kernel evaluations. We also use the identified value of $t_{h}$ in the reference approach described in Section 6.3. We use Matlab’s ode45 with absolute and relative tolerances of $10^{-12}$ to simulate the approximate system of ODEs (61), and we use $K_{x}=24\cdot 10^{3}$ quadrature points to evaluate the state error in (104), which corresponds to a time step size of $\Delta t=10^{-3}$ mo. Furthermore, we use $K_{\alpha}=10^{6}$ quadrature points to evaluate the kernel error. For the results obtained with the numerical methods described in Appendix B, we use a memory horizon of $\Delta t_{h}=t_{f}-t_{0}=24$ mo and the above time step size, and for the method for stiff DDEs, we use a function tolerance of $10^{-12}$ and an optimality tolerance of $10^{-8}$ when we solve the involved residual equations with Matlab’s fsolve.
@@
The results demonstrate that for this example, the Erlang mixture approximation obtained with the proposed least-squares approach is several orders of magnitude more accurate than the reference approach based on the theoretical expressions for the coefficients. In both cases, there is a high correlation between the error of the kernel and the error of the approximate state trajectory. The kernel and state error is higher for $M=32$ than $16$ because the objective function in the NLP (95) uses a coarser resolution in the rectangle rule than the error in (104). Finally, the state error decreases quadratically with the time step size for the two numerical methods described in Appendix B because the involved integral have been discretized with first-order methods and the error is squared in (104).
@@
Figure 4: The state (top left) and kernel (bottom left) errors for Erlang mixture approximations of different orders, $M$, obtained with the proposed least-squares approach and the reference approach based on the theoretical expressions for the coefficients. The bottom right shows the state and kernel errors against each other, and the top right plot shows the state error obtained with the numerical approaches for non-stiff and stiff DDEs for different time step sizes, $\Delta t$. The black solid line is proportional to $\Delta t^{2}$.
@@ 8.1 Modified logistic equation
We use the parameter values in Table 1, and the initial state is $x(t)=0.9$ for $t\leq t_{0}$. We use the same parameters as in Section 7 to identify Erlang mixture kernels of order $M=100$ in the bifurcation analysis with respect to the logistic growth rate, $\sigma$, and $M=350$ in the analysis with respect to the kernel parameter $\mu_{2}$. In Figure 5, the first row shows the eigenvalues of the Jacobian of the right-hand side function in the approximate system of ODEs (61). For completeness, we consider a very large range of values for $\sigma$, but for numerical reasons, we have not been able to investigate larger values of $\mu_{2}$. The second row shows the largest real part of the eigenvalues and when the parameters become sufficiently large, the steady state $\bar{x}=\kappa$ becomes unstable. The two bottom rows show simulations obtained with the numerical method for non-stiff DDEs for the parameter values indicated in the second row (the colors match). We use a memory horizon of $\Delta t_{h}=t_{f}-t_{0}=24$ mo and $10,000$ time steps. In both cases, the stability analysis based on the approximate set of ODEs (61) accurately predicts when the steady state is stable, marginally stable (where the largest real part of the eigenvalues is equal to zero), and unstable. For large values of $\sigma$, a limit cycle appears whereas for large values of $\mu_{2}$, the population density oscillates unstably. Figure 6 shows the identified kernels, the corresponding error, and the coefficients for different values of $\mu_{2}$. There is a similarity between the coefficients and the shape of the approximated kernel, e.g., in terms of bimodality. Furthermore, as $\mu_{2}$ increases, more sharp spikes occur in the coefficients.
@@ 8.2 Nuclear fission
The dilution rate, $D\in\mathbb{R}_{\geq 0}$, is the ratio between the volumetric inlet and outlet flow rate and the reactor core volume, $\kappa\in\mathbb{R}_{\geq 0}$ is a proportionality constant, and $H\in\mathbb{R}_{\geq 0}$ is the ratio between the power production proportionality constant and the heat capacity of the reactor core. Furthermore, the production rate $R:\mathbb{R}\rightarrow\mathbb{R}^{n}$ is defined in terms of the stochiometric matrix $S\in\mathbb{R}^{N_{r}\times n}$ where $N_{r}=n\in\mathbb{N}$ is the number of reactions and $r:\mathbb{R}\rightarrow\mathbb{R}_{\geq 0}^{N_{r}}$ is a vector of reaction rates:
@@ 9 Conclusions
In this paper, we propose an approximate approach for simulating and analyzing DDEs with distributed time delays based on conventional methods for ODEs. Specifically, we approximate the involved kernel by an Erlang mixture kernel and transform the resulting DDEs to ODEs using the LCT. Furthermore, we prove that if the kernel is continuous and bounded, an Erlang mixture kernel approximation of infinite order converges to the true kernel as the (common) rate parameter converges to infinity. Next, we compare the steady states of the original DDEs and the approximate system of ODEs as well as the corresponding stability criteria. Additionally, we propose an approach for determining optimal Erlang mixture approximations based on bisection and least-squares estimation, and we use a numerical example to demonstrate that its accuracy and rate of convergence can be higher than using the theoretical expressions for the coefficients obtained from the proof of convergence. Finally, we present two numerical examples that demonstrate the efficacy of the proposed approach, and we compare the results with those obtained by two numerical methods (for non-stiff and stiff systems) formulated directly for DDEs with distributed time delays.
@@ Appendix A Derivation of the linear chain trick
(113a) ˙$\dot{\ell}_{0,i}(t)$ $=-a_{i}\ell_{0,i}(t),$ $i$ $=1,\ldots,n_{z},$
(113b) ˙$\dot{\ell}_{m,i}(t)$ $=a_{i}(\ell_{m-1,i}(t)-\ell_{m,i}(t)),$ $m$ $=1,\ldots,M_{i},$ $i$ $=1,\ldots,n_{z},$
@@
In order to derive differential equations for the auxiliary memory states, we use Leibniz’ integral rule [39, Thm. 3, Chap. 8] to differentiate (111):
@@
(116) $\ell_{m,i}(0)$ =$=\begin{cases}a_{i},&\text{for}~{}m=0,\\ 0,&\text{for}~{}m=1,\ldots,M_{i},\end{cases}$ $i$ $=1,\ldots,n_{z},$

### v2

@@ 6.2 Identify kernel parameters
where the quadrature points are $t_{k}=k\Delta t$ for $k=0,\ldots,N-1$, $\Delta t\in\mathbb{R}_{>0}$ is the width of the rectangles, and $a_{\min}\in\mathbb{R}_{>0}$ is an arbitrary lower bound, which ensures that only positive values of $a$ are feasible. Furthermore, the objective function involves the regular kernel, $\alpha:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}$, and the $M$’th order Erlang mixture kernel, $\hat{\alpha}:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}$, with coefficients $\{c_{m}\}_{m=0}^{M}$ and rate parameter $a$, evaluated at the quadrature points.
@@ 6.4 Implementation
We implement the evaluation of the objective function in the NLP (120) and its first- and second-order derivatives in C, and we use Matlab’s MEX interface to call the C routines from Matlab. Furthermore, we use Matlab’s fmincon to approximate the solution to the NLP, and we provide analytical first- and second-order derivatives. We exploit that the recursion $\ell_{m}(t)=(at/m)\ell_{m-1}(t)$ for $m=1,\ldots,M$ where $\ell_{0}(t)=ae^{-at}$ is less prone to rounding errors than a straightforward evaluation of the expression (11) in the definition, and we use either double or long double precision when evaluating the Erlang kernels, the Erlang mixture kernels, and their derivatives. We simulate the approximate system of ODEs (73) using either Matlab’s ode45 or ode15s, and depending on the order of the Erlang mixture kernel, we implement the matrices $A$, $B$, and $C$ as either dense or sparse. Furthermore, we provide the analytical Jacobian of the right-hand side function of the approximate system of ODEs (73) to the simulator. We implement the numerical method for non-stiff DDEs described in Appendix B.1 in both Matlab and C (with a MEX interface). The numerical method from Appendix B.2 is implemented in Matlab, we use fsolve to solve the residual equations, and we provide the analytical Jacobian. In all three implementations, the right-hand side function, $f$, the delay function, $h$, and the kernel, $\alpha$, are evaluated in Matlab. Finally, in Section 8, we carry out parallel simulations and kernel approximations using a high-performance computing cluster [14].
@@ 7 Numerical test of the accuracy
In this section, we construct a DDE in the form (2b)–(3) with a known solution and use it to investigate the accuracy of the Erlang mixture approximation and the numerical solution of the approximate system of ODEs (73). Furthermore, we assess the convergence rate of the two numerical methods for solving initial value problems in the form (2)–(3) described in Appendix B. Finally, we compare the accuracy of the Erlang mixture approximation with that of one of the exponential mixture approximations described in [18] for a system of DDEs involving a gamma kernel, which describes chemotherapy-induced myelosuppression [33].
7.1 Modified logistic equation
@@
(129) $E_{x}$ $=\sum_{n=0}^{N_{x}-1}(\hat{x}(t_{n+1})-x^{*}(t_{n+1}))^{2}\Delta t,$ $E_{\alpha}$ $=\sum_{k=0}^{N_{\alpha}-1}(\hat{\alpha}(t_{k})-\alpha(t_{k}))^{2}\Delta t.$
@@
Here, $\hat{x}$ and $\hat{\alpha}$ denote the approximations. Note that the kernel error, $E_{\alpha}$, is analogous to the objective function in (120a) but may contain a different number of terms. We have used the model parameter values $\sigma=4$ mo-1 and $\kappa=1$ and the dilation parameter $\gamma=10$ mo, and the initial and final times are $t_{0}=0$ mo and $t_{f}=24$ mo. When we determine the Erlang mixture approximation of a given order, $M$, we use $N=10^{3}$ points in the objective function in (120a), and we use $a_{\min}=1$ mo-1. Furthermore, we use a threshold of $\epsilon=10^{-14}$ and a convergence tolerance of $10^{-15}$ to determine the approximation interval length, $t_{h}$. Finally, we solve the NLP (120) using Matlab’s fmincon with an optimality tolerance of $10^{-10}$ and a step size tolerance below machine precision, and we use double precision in the involved kernel evaluations. We also use the identified value of $t_{h}$ in the reference approach described in Section 6.3. We use Matlab’s ode45 with absolute and relative tolerances of $10^{-12}$ to simulate the approximate system of ODEs (73), and we use $N_{x}=24\cdot 10^{3}$ quadrature points to evaluate the state error in (129), which corresponds to a time step size of $\Delta t=10^{-3}$ mo. Furthermore, we use $N_{\alpha}=10^{6}$ quadrature points to evaluate the kernel error. For the results obtained with the numerical methods described in Appendix B, we use a memory horizon of $\Delta t_{h}=t_{f}-t_{0}=24$ mo and the above time step size, and for the method for stiff DDEs, we use a function tolerance of $10^{-12}$ and an optimality tolerance of $10^{-8}$ when we solve the involved residual equations with Matlab’s fsolve.
@@
The results demonstrate that for this example, the Erlang mixture approximation obtained with the proposed least-squares approach is several orders of magnitude more accurate than the reference approach based on the theoretical expressions for the coefficients. In both cases, there is a high correlation between the error of the kernel and the error of the approximate state trajectory. For the least-squares approach, the convergence rates are faster than polynomial, and for the reference approach, the convergence rates are quadratic for sufficiently large values of $M$ (see also Figure 9, and note that the errors are squared in (129)). The kernel and state error is higher for $M$ equal to 64 and 128 than for $32$ because the objective function in the NLP (120) uses a coarser resolution in the rectangle rule than the error in (129). Finally, the state error decreases quadratically with the time step size for the two numerical methods described in Appendix B because the involved integral and differential equation have been discretized with first-order methods.
@@
Figure 3: Convergence analysis for the modified logistic equation. The state (top left) and kernel (bottom left) errors for Erlang mixture approximations of different orders, $M$, obtained with the proposed least-squares approach and the reference approach based on the theoretical expressions for the coefficients. The bottom right plot shows the state and kernel errors against each other, and the top right plot shows the state error obtained with the numerical approaches for non-stiff and stiff DDEs for different time step sizes, $\Delta t$. The black solid lines in the left column are proportional to $1/(M+1)^{2}$, and the one in the top right is proportional to $\Delta t^{2}$.
7.2 Myelosuppression
In this section, we compare the accuracy of the Erlang mixture approximation proposed in this work with an exponential mixture approximation (see Section F) used by Guglielmi and Hairer [18] for a system of DDEs describing myelosuppression. The system describes the evolution of precursor cells, $y:\mathbb{R}\rightarrow\mathbb{R}_{\geq 0}$, which ultimately form granulocytes, $w:\mathbb{R}\rightarrow\mathbb{R}_{\geq 0}$:
(130a) $\dot{y}(t)$ $=\left(\kappa\left(\frac{w_{0}}{w(t)}\right)^{\gamma}-k_{s}C(t)-\kappa\right)y(t),$
(130b) $\dot{w}(t)$ $=-\kappa w(t)+\kappa\int_{-\infty}^{t}\alpha(t-s)y(s)\,\mathrm{d}s,$
(130c) $\dot{C}(t)$ $=-\frac{V_{\max}C(t)}{K_{m}+C(t)}.$
The purpose of the model is to study the effect of a bolus injection of a drug (fluoracil) and $C:\mathbb{R}\rightarrow\mathbb{R}_{\geq 0}$ is the concentration of the drug in the plasma, which is described by the Michaelis-Menten model in (130c). The constants are $\kappa=(1-\lambda)/55.6$ h-1, $\lambda=-0.46$, $w_{0}=14.4\cdot 10^{9}$ cells/L, $\gamma=0.507$, $k_{s}=0.0213$ L/(mg h), $V_{\max}=100$ mg/(L h), and $K_{m}=22$ mg/L, and the gamma kernel is given by
(131) $\alpha(t)$ $=\frac{\kappa^{1-\lambda}}{\Gamma(1-\lambda)}t^{-\lambda}e^{-\kappa t},$
where $\Gamma:\mathbb{C}\rightarrow\mathbb{C}$ is the gamma function. The initial and final times are $t_{0}=0$ h and $t_{f}=700$ h, and the initial states are $y(t)=w_{0}$ for $t\in(-\infty,t_{0}]$, $w(t_{0})=w_{0}$, and $C(t_{0})=A_{0}/V$ where $A_{0}=127$ mg/kg and $V=1.03$ L/kg. We use a threshold of $\epsilon=10^{-3}$ in the computation of $t_{h}$, and we use $N=10^{4}$ points in the objective function. The initial guess of the rate parameter, $a$, is $\kappa$, and we use $a_{\min}=10^{-3}$ h-1. As the system is stiff, we use ode15s to simulate the approximate system of ODEs (73), and we compare to the error of the exponential mixture approximation described in [18], which depends on the tolerance $\epsilon_{r}\in\mathbb{R}_{>0}$ that also indirectly determines the number of terms in the approximation. Lower tolerances lead to more terms, and we consider $\epsilon_{r}\in\{10^{-1},10^{-3},10^{-5},10^{-7},10^{-9},10^{-11}\}$. Furthermore, we compare the state obtained with the Erlang mixture approach to that obtained with $\epsilon_{r}=10^{-11}$ using $N_{x}=700\cdot 10^{3}$ points in time, corresponding to a time step size of $\Delta t=10^{-3}$ h. The remaining parameters are the same as in Section 7.1.
The results are shown in Figure 4, and as for the example in Section 7.1, the least-squares approach described in Section 6 is significantly more accurate than the reference approach based on the theoretical expressions for the coefficients. However, the error is generally higher than for the modified logistic equation (see Figure 3), the convergence rates are no longer faster than polynomial for the least-squares approach, and they are no longer quadratic for the reference approach based on the theoretical expressions for the coefficients. The figure also shows that Guglielmi and Hairer’s approach is significantly more accurate (note the difference in the vertical axes). The error of the Erlang mixture approximation is primarily high for $t$ close to zero. The reason for this may be that the derivative of the true kernel has a singularity at zero whereas for finite values of $M$ and $a$, the derivative of the Erlang mixture approximation does not. This may also describe why the convergence rates of the Erlang mixture approximation are reduced. However, we emphasize that the Erlang mixture approximation can be used for a broader class of kernels (bounded and continuous) than the exponential mixture approximations.
Figure 4: Convergence analysis for the myelosuppression model. The state (top left) and kernel (bottom left) errors for Erlang mixture approximations of different orders, $M$, obtained with the least-squares approach and the reference approach based on theoretical values of the coefficients. The kernel error is shown for Guglielmi and Hairer’s approach in the bottom right, and the true and approximate kernels are shown in the top right for $M=128$ and $\epsilon_{r}=10^{-11}$ (corresponding to 309 terms). The black solid line in the top left is proportional to $1/(M+1)^{2}$ whereas the black dashed line in the bottom left is proportional to $1/(M+1)^{3/2}$.
@@ 8.1 Modified logistic equation
We use the parameter values in Table 1, and the initial state is $x(t)=0.9$ for $t\leq t_{0}$. We use the same parameters as in Section 7.1 to identify Erlang mixture kernels of order $M=100$ in the bifurcation analysis with respect to the logistic growth rate, $\sigma$, and $M=350$ in the analysis with respect to the kernel parameter $\mu_{2}$. In Figure 5, the first row shows the eigenvalues of the Jacobian of the right-hand side function in the approximate system of ODEs (73). For completeness, we consider a very large range of values for $\sigma$, but for numerical reasons, we have not been able to investigate larger values of $\mu_{2}$. The second row shows the largest real part of the eigenvalues and when the parameters become sufficiently large, the steady state $\bar{x}=\kappa$ becomes unstable. The two bottom rows show simulations obtained with the numerical method for non-stiff DDEs for the parameter values indicated in the second row (the colors match). We use a memory horizon of $\Delta t_{h}=t_{f}-t_{0}=24$ mo and $10,000$ time steps. In both cases, the stability analysis based on the approximate set of ODEs (73) accurately predicts when the steady state is stable, marginally stable (where the largest real part of the eigenvalues is equal to zero), and unstable. For large values of $\sigma$, a limit cycle appears whereas for large values of $\mu_{2}$, the population density oscillates unstably. Figure 6 shows the identified kernels, the corresponding error, and the coefficients for different values of $\mu_{2}$. There is a similarity between the coefficients and the shape of the approximated kernel, e.g., in terms of bimodality. Furthermore, as $\mu_{2}$ increases, more sharp spikes occur in the coefficients.
@@ 8.2 Nuclear fission
The dilution rate, $D\in\mathbb{R}_{\geq 0}$, is the ratio between the volumetric inlet and outlet flow rate and the reactor core volume, $\kappa\in\mathbb{R}_{\geq 0}$ is a proportionality constant, and $H\in\mathbb{R}_{\geq 0}$ is the ratio between the power production proportionality constant and the heat capacity of the reactor core. Furthermore, the production rate $R:\mathbb{R}\rightarrow\mathbb{R}^{n}$ is defined in terms of the stoichiometric matrix $S\in\mathbb{R}^{N_{r}\times n}$ where $N_{r}=n\in\mathbb{N}$ is the number of reactions and $r:\mathbb{R}\rightarrow\mathbb{R}_{\geq 0}^{N_{r}}$ is a vector of reaction rates:
@@ 9 Conclusions
In this paper, we propose an approximate approach for simulating and analyzing DDEs with distributed time delays based on conventional methods for ODEs. Specifically, we approximate the involved kernel by an Erlang mixture kernel and transform the resulting DDEs to ODEs using the LCT (Theorem 5.1). We prove that if the kernel $\alpha$ is regular (Definition 3.1) and the coefficients are chosen as in (17), then an infinite-order Erlang mixture approximation $\hat{\alpha}$ (Definition 4.2 where $M$ is infinite) converges pointwise to $\alpha$ (Theorem 4.26) and the integral of the absolute error converges to zero (Theorem 4.37) as the rate parameter, $a$, goes to infinity. Furthermore, we show that accurate stability analyses of the steady states of the original system of DDEs (2b)–(3) can be performed based on the approximate system of ODEs (73) if the kernel $\alpha$ is also exponentially bounded (Theorem 5.14). We also show that the solution to the approximate ODEs converges to that of the original system (Theorem 5.22). Additionally, we propose an approach for determining optimal Erlang mixture approximations based on bisection and least-squares estimation, and we use numerical examples to demonstrate that its accuracy and rate of convergence can be higher than using the theoretical expressions for the coefficients in (17). We also compare with the accuracy of an exponential mixture approximation from the literature. Finally, we present two numerical examples that demonstrate the efficacy of the proposed approach, and we compare the results with those obtained by two numerical methods (for non-stiff and stiff systems) formulated directly for DDEs with distributed time delays.
Future work may involve the derivation of error bounds, analysis of the error from truncating infinite-order Erlang mixture kernels (as in Section 6), an extension to Erlang mixture approximations with multiple rate parameters, and an extension to systems with time-varying kernels.
@@ Appendix A Derivation of the linear chain trick
(140a) $\dot{\ell}_{0,i}(t)$ $=-a_{i}\ell_{0,i}(t),$ $i$ $=1,\ldots,n_{z},$
(140b) $\dot{\ell}_{m,i}(t)$ $=a_{i}(\ell_{m-1,i}(t)-\ell_{m,i}(t)),$ $m$ $=1,\ldots,M_{i},$ $i$ $=1,\ldots,n_{z},$
@@
In order to derive differential equations for the auxiliary memory states, we use Leibniz’ integral rule [42, Thm. 3, Chap. 8] to differentiate (138):
@@
(143) $\ell_{m,i}(0)$ $=\begin{cases}a_{i},&\text{for}~m=0,\\ 0,&\text{for}~m=1,\ldots,M_{i},\end{cases}$ $i$ $=1,\ldots,n_{z},$